OpenStudy (anonymous):

find the equation of the quadratic function with roots -6 and -4 and through the point (4,40) ...medal provided

OpenStudy (imstuck):

To find the equation, multiply the roots, which are (x+6)(x+4). That comes to\[x ^{2}+10x+24\]

OpenStudy (imstuck):

Now we have to worry about the point (4,40).

OpenStudy (imstuck):

Line of symmetry is x = -5...

OpenStudy (zzr0ck3r):

a(x+6)(x+4)=y plug in (-6,-4) solve for a

OpenStudy (imstuck):

why plug in (-6,-4)?

OpenStudy (imstuck):

Do you mean a(4+6)(4+4)=40 and solve for a?

OpenStudy (imstuck):

The point (4,40) needs to go in there, doesn't it?

OpenStudy (imstuck):

I get that a = .5

OpenStudy (zzr0ck3r):

yes I did

OpenStudy (zzr0ck3r):

sorry

OpenStudy (imstuck):

So it is \[.5(x ^{2}+10x+24)\]

OpenStudy (zzr0ck3r):

or as mathematicians say (1/2)(x^2+10x+24)

OpenStudy (zzr0ck3r):

:P

OpenStudy (zzr0ck3r):

are you an engineer @IMStuck ?

OpenStudy (imstuck):

Which gives us:\[.5x ^{2}+10(.5)x+24\]which is\[\frac{ 1 }{ 2 }x ^{2}+5x+24=y\]

OpenStudy (imstuck):

No...not an engineer...high school math is my thing! Teacher, dontcha know?!

OpenStudy (imstuck):

You're way way smarter than I am, @zzr0ck3r !!!!!

OpenStudy (zzr0ck3r):

engineers love decimals

OpenStudy (zzr0ck3r):

lol that has nothing to do with it

OpenStudy (zzr0ck3r):

and I doubt it.

OpenStudy (anonymous):

@IMStuck how do you know what the line of symmetry is?

OpenStudy (imstuck):

Well the line of symmetry is exactly halfway between the x intercepts. Since this is an x^2 parabola, the graph opens upward and the x point halfway between -6 and -4 is -5.

OpenStudy (anonymous):

@IMStuck ohh.. like the equation x+x/ 2 ...

OpenStudy (imstuck):



OpenStudy (anonymous):

@IMStuck so then howd you get .5(x2+10x+24)